### v3

# slow-fast system and dynamics near infinity

Abstract.
We focus on the qualitative analysis of the phase portraits arising in the three-parameter FitzHugh-Nagumo system and its compactified form. The investigation is split into three parameter-dependent cases. In one of these cases, the system displays a double-zero bifurcation with $\mathbb{Z}_{2}$–symmetry, a singularity of codimension two. For this case, we provide explicit descriptions of the bifurcation/transition curves unfolding the singularity, including Pitchfork, Hopf, Belyakov, and Double Homoclinic. Further, we present the corresponding bifurcation diagrams.
We bridge this analysis with the theory on the framework of slow-fast family and to the presence of canards.
We complete our study with the analysis of the dynamics “near infinity” for the family under consideration. This study complements the work summarised in Georgescu, Rocşoreanu and Giurgiţeanu, Global Bifurcations in FitzHugh-Nagumo Model, Trends in Mathematics: Bifurcations, Symmetry and Patterns (2003).
Alexandre A. P. Rodrigues1,† $\dagger$}Corresponding author and Nasrin Sadri2
1 Lisbon School of Economics and Management, Lisbon University, Centro de Matemática Aplicada à Previsão e Decisão Económica,
Rua do Quelhas 6, 1200-781 Lisboa, Portugal,
2 School of Mathematics, Statistics and Computer Science, College of Science, University of Tehran, Tehran, Iran and School of Mathematics, Institute for Research in Fundamental Sciences (IPM), P. O. Box 19395-5746, Tehran, Iran
2020 Mathematics Subject Classification:
Primary: 34C60. Secondary: 34E17, 34E15, 37G15, 37G35
1. Introduction
Information in nerve fibers is encoded through action potentials (electrical membrane changes). A. Hodgkin and A. Huxley, in 1939, succeeded in recording an action potential from within a neuron for the first time, using microelectrodes inserted into the squid’s giant axon. In their experiments, they demonstrated that the action potentials arise from two effects:
  • a fast inward current mediated by sodium (Na+) ions, and
  • a slow developing outward current produced by potassium (K+) ions.
Hodgkin and Huxley showed that the membrane’s permeabilities to $\text{Na}^{+}$ and $\text{K}^{+}$ are controlled separately, with their conductances varying in time and as functions of the membrane voltage. They formulated a quantitative description of the action potential in a single neuron, now referred to as the Hodgkin-Huxley (HH) model [11]. Details and an overview may be found in the review [3].
1.1. From the Hodgkin-Huxley model to the FitzHugh-Nagumo equations
Hodgkin and Huxley represented the observed current variations using probabilistic variables that describe the closing and opening of ion channels. The Hodgkin-Huxley framework characterises the electrical dynamics of these membrane channels by modelling the flow of $\text{Na}^{+}$ and $\text{K}^{+}$ ions across the membrane. In the HH framework, the dynamics are governed by four ODEs, each one associated with a specific state variable – [11, Eq. 1]. This system is nonlinear, making it analytically difficult to solve explicitly. Nonetheless, numerical simulations allow investigation of specific features and general dynamics, including the presence of excitability and oscillations.
The ability of the HH model to accurately reproduce these excitable dynamics represented a key insight, offering a complete explanation of the recorded action potential and the mechanisms underlying neural excitability (cf. [11]). While the HH model successfully reproduces many neuronal physiological phenomena, it is very complex. In [7], Richard FitzHugh introduced a simplified model to capture the dynamics of neuronal excitability, which was refined by Jinichi Nagumo two years later [19]. This system is known as the FitzHugh-Nagumo (FHN) model. FitzHugh concentrated the study on preserving the key dynamical characteristics of the HH model, specifically and oscillatory and excitability behaviour.
1.2. The FitzHugh-Nagumo equations
FitzHugh started with the oscillator equation (introduced in 1920) by Balthasar Van der Pol [25], which supports oscillations resembling relaxation-like dynamics. This means there are periods of “low” and “high” states characterised through fast transitions between them.
The Van der Pol (VdP) equation is derived from the basic differential equation describing a damped harmonic oscillator. He replaced the constant damping by a damping function that has a quadratic dependence on $x,$ resulting in the nonlinear differential equation ($\dot{x}$ denotes the first derivative of the physical position $x$ with respect to $t$ and $c\in\mathbb{R}^{+}_{0}$):
$\ddot{x}+c(x^{2}-1)\dot{x}+x=0.$
The system exhibits “effective” damping only when $|x|<1$; for $x^{2}>1$, the nonlinear term leads to amplification. To analyse the dynamics of the VdP equation, the Liénard map can be employed [16]
$y=\frac{\dot{x}}{c}+\left(\frac{x^{3}}{3}-x\right),$
giving rise to a system of two differential equations (on the plane):
$\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}x,\end{array}\right$. (1.4)
from which the time-scale separation of the equations becomes evident.
For $c\gg 1$, the first variable changes rapidly of the order of $\mathcal{O}(c)$, whereas the second evolves much more slowly, $\mathcal{O}(1/c),$ where $\mathcal{O}$ denotes the standard Landau notation (given a real map $f:\mathbb{R}\to\mathbb{R}$ and $c\in\mathbb{R}$, we say that $f$ is of order $\mathcal{O}(c)$ if there exist $M\in\mathbb{R}^{+}$ and $x_{0}\in\mathbb{R}$ such that $|f(x)|<Mc$, for $x>x_{0}$). Based on the VdP oscillator, FitzHugh considered the equation:
$\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)+z\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}(x-a+by)\end{array}\right$. (1.5)
where (see [3] and Remark 1.1):
  1. (1)
    the parameter $z$ mimics the membrane current density;
  2. (2)
    the variable $x$ corresponds to the membrane voltage and the $\text{Na}^{+}$ activation and
  3. (3)
    the variable $y$ is related to the $\text{Na}^{+}$ inactivation and the $\text{K}^{+}$ activation.
Equations (1.5) act as an activator-inhibitor framework. Arimoto, Nagumo, and Yoshizawa proved that it can be realized as an electrical circuit [19]. Although the FHN model is primarily used to describe neuronal and cardiac systems, it can also be applied in a range of other biological contexts.
The FHN equations provide a simplified framework for representing interconnected negative and positive feedback loops, enabling the generation of a wide range of responses such as switches, pulses, and stable oscillations [24]. Within the wide range of models used to investigate cardiac cells, the FHN model is notable for being one of the simplest and most extensively researched, effectively representing the general dynamic characteristics of cardiac cells [1, 4].
Remark 1.1.
System (1.5) mimics an excitable system and there are limitations regarding the interpretation of the original variables of the HH model. In the rest of the paper, we will take a cautious approach and we stay within a mathematical motivation grounded in bifurcations and dynamical systems.
1.3. Novelty
Results in [21, pp. 180] synthesised the global bifurcation diagram for the FHN model. The result comes from integrating all the local bifurcation diagrams obtained in Chapters 2–4, as if fitting together a “huge puzzle”. See Section 2.4 of [20].
One of the novelties of this paper is the dissection of the bifurcations diagrams of [21, pp. 180] and provide a more complete bifurcation analysis of model (1.5), namely the location of Hopf, Pitchfork and Double-zero bifurcations, which will be specified in Subsection 1.4. Besides the case-study $c=\mathcal{O}(1)$, we also study the asymptotic case
$c\to\pm\infty\qquad(\Leftrightarrow\quad\varepsilon=1/c^{2}\to 0),$
where canards are observed [12, 14], and we finish the analysis with the compactification of the phase portrait of (1.5) on the Poincaré disc.
To the best of our knowledge, this procedure applied to equation (1.5) is new, providing additional information about the trajectories which come from or tend to infinity. We classify topologically all phase portraits of system (3.1) “near infinity”, represented on the Poincaré disc (i.e. within the compactification of $\mathbb{R}^{2}$ by including the circle at infinity, $\mathbb{S}^{1}$). Note that we do not claim to have derived complete phase portraits.
1.4. Structure
This paper is structured as follows: Section 2 introduces the definitions and preliminary concepts of quasi-homogeneous blow-up and Poincaré compactification. In Section 3, we discuss the finite equilibria of (3.1) for the following three cases:
Case A:
$a=0$
Case B:
$b=0$
Case C:
$a\neq 0,\,0<b<1$
and investigate possible bifurcations, including Pitchfork, Double Homoclinic and Hopf bifurcations. The general bifurcation analysis of (1.5) is difficult to tackle; however we have been able to make some progress in Cases A, B and C. In Section 4, we analyse the asymptotic dynamics of (1.5) when $c\to\pm\infty$ and we relate the existence of canards with the periodic solution emerging from the Hopf bifurcation.
Section 5 is dedicated to studying the dynamics “near infinity” for (1.5) on the Poincaré disc. The phase portraits for each connected component in Cases A, B and C are illustrated. This complements the work started in [8]. Section 6 finishes the paper.
In this paper, we aim to provide a self-contained exposition, integrating all topics pertinent to the proofs. Illustrative figures have been included for clarity, and all results are supported by numerical simulations performed using Matlab R2015b and pplane9.
2. Preparatory section
This section provides definitions for polynomial vector fields on $\mathbb{R}^{2}$ that will be used in the sections that follow. Let $f_{(a,b,c)}$ be a smooth vector field on $\mathbb{R}^{2}$ whose flow is given by the unique solution $X(t)=\varphi(t,X)\in\mathbb{R}^{2}$ of the three-parameter family
$\dot{X}=f_{(a,b,c)}(X),\qquad X(0)=X_{0}\in\mathbb{R}^{2},$ (2.1)
[…]
2.1. Useful terminology
The center manifold of a non-hyperbolic equilibrium $p\in\mathbb{R}^{2}$ of (2.1) is the set of solutions whose behaviour around $p$ is controlled neither by the exponential attraction of the stable manifold nor by the exponential repulsion of the unstable manifold.
If $Df_{(a,b,c)}$ evaluated at an equilibrium $p$, $Df_{(a,b,c)}(p)$, has an eigenvalue with zero real part, the center manifold plays an important role and this is the set where bifurcations might occur.
[…]
Throughout this paper, we study the Double-zero (DZ) singularity with symmetry $\mathbb{Z}_{2}$ for a family of differential equations corresponding to the case described in [27, pp. 400]. The unfolding of this singularity of codimension two involves lines of Belyakov transitions, Pitchfork, Hopf, Double Homoclinic and saddle-node bifurcation of limit cycles.
A Hopf bifurcation is termed supercritical when the first Lyapunov coefficient of its normal form is negative, leading to the emergence of a stable small-amplitude limit cycle, and subcritical when this coefficient is positive, in which an unstable small-amplitude limit cycle is created [26]. For full details on these bifurcations and the criteria that guarantee their appearance, see [15, 18, 27].
Remark 2.1.
In this paper, we refer to an equilibrium point $p$ as undergoing a Belyakov transition if at least one pair of eigenvalues of $Df_{(a,b,c)}(p)$ changes from real to complex conjugate or vice versa, while the sign of their real part remains unchanged. Although such a transition is typically considered in higher-dimensional systems, we adopt this terminology here for the sake of brevity and clarity.
2.2. Quasi-homogeneous blow-up method
Quasi-homogeneous blow-ups are used to analyze and simplify the behaviour of differential equations near non-hyperbolic singularities by rescaling the variables according to their natural weights. This technique transforms the singularity into a more regular, analyzable object. Although homogeneous blow-ups
$\phi:\mathbb{S}^{1}\times\mathbb{R}\to\mathbb{R}^{2},\quad(\theta,\rho)\mapsto\left(\rho\cos\theta,\,\rho\sin\theta\right),$
suffice for the study of isolated singularities of an analytic vector field, the quasi-homogeneous approach (scaling the variables with different degrees) is considerably more efficient. Here we review this method briefly – for more details and examples, see [5, Section 3.3].
Consider the polynomial vector field $f=(P,Q)$ associated with the planar differential equation:
$\left\{\begin{array}[]{l}\dot{x}_{1}=P(x_{1},x_{2})\\ \\ \dot{x}_{2}=Q(x_{1},x_{2}).\end{array}\right$. (2.2)
In the language of singularity theory, we have:
$f(x_{1},x_{2})=P(x_{1},x_{2})\frac{\partial}{\partial x_{1}}+Q(x_{1},x_{2})\frac{\partial}{\partial x_{2}},$
with an isolated singular point at the origin $(0,0)$. Since $P$ and $Q$ are polynomial we may find $p_{ij},q_{ij}\in\mathbb{R}$ such that:
$P(x_{1},x_{2})=\sum_{i+j\geq 1}p_{ij}x_{1}^{i}x_{2}^{j}\qquad\text{and}\qquad Q(x_{1},x_{2})=\sum_{i+j\geq 1}q_{ij}x_{1}^{i}x_{2}^{j}.$
Definition 2.2.
The quasi-homogeneous blow-up is defined by
$\phi:\mathbb{S}^{1}\times\mathbb{R}\to\mathbb{R}^{2},\quad(\theta,\rho)\mapsto\left(\rho^{\alpha}\cos\theta,\,\rho^{\beta}\sin\theta\right)=(x_{1},x_{2}),$
for some $\alpha,\beta\in\mathbb{N}.$
If $(\alpha,\beta)=(1,1)$, this corresponds to the standard homogeneous case. Doing the quasi-homogeneous blow-up (of the previous definition) on $f=(P,Q)$, it yields the following differential equation:
$\dot{\rho}=\frac{P\rho^{\beta}\cos\theta+Q\rho^{\alpha}\sin\theta}{\rho^{\alpha+\beta-1}(\alpha\cos^{2}\theta+\beta\sin^{2}\theta)},\qquad\dot{\theta}=\frac{\alpha Q\rho^{\alpha}\cos\theta-\beta P\rho^{\beta}\sin\theta}{\rho^{\alpha+\beta}(\alpha\cos^{2}\theta+\beta\sin^{2}\theta)},$
in which $P$ and $Q$ are computed at $(\rho^{\alpha}\cos\theta,\rho^{\beta}\sin\theta)$. A common factor $\rho^{d}>0$ can be cancelled for a suitable choice of $d>0$.
Throughout this article, we will be interested in quasi-homogeneous directional blow-ups in a given direction. This technique isolates and desingularizes the dynamics of trajectories approaching the singularity along that direction, creating a chart where the vector field becomes simpler and its behaviour can be analyzed separately from other directions.
Definition 2.3.
The quasi-homogeneous directional blow-up along the positive (resp. negative) $x$ axis is given by the map:
$\overline{f}^{x}_{+}:\quad(\bar{x},\bar{y})\mapsto(\bar{x}^{\alpha},\bar{x}^{\beta}\bar{y}),$
$\overline{f}^{x}_{-}:\quad(\bar{x},\bar{y})\mapsto(-\bar{x}^{\alpha},\bar{x}^{\beta}\bar{y}),$
and the quasi-homogeneous directional blow-up along the positive (resp. negative) $y$ axis is described by:
$\overline{f}^{y}_{+}:\quad(\bar{x},\bar{y})\mapsto(\bar{x}\bar{y}^{\alpha},\bar{y}^{\beta}),$
$\overline{f}^{y}_{-}:\quad(\bar{x},\bar{y})\mapsto(\bar{x}\bar{y}^{\alpha},-\bar{y}^{\beta}),$
where $\bar{x}$ and $\bar{y}$ denote new variables and $\alpha,\beta\in\mathbb{N}.$
Among all possible pairs $(\alpha,\beta)\in\mathbb{N}\times\mathbb{N}$ of the definition of quasi-homogeneous blow-ups, a natural question is which choice is the most “efficient” (in the sense that we desingularize the equilibria in the minimum number of steps) for performing a quasi-homogeneous blow-up to simplify and resolve a singularity. Selecting such an efficient pair balances the relative importance of the variables, making the leading-order dynamics near the singularity clear. The Newton polygon provides a systematic tool to guide this choice, which we now introduce (adapted from [5]).
Definition 2.4.
We define the support of $f$ as the subset
$S=\{(m-1,n)\mid p_{mn}\neq 0\}\cup\{(m,n-1)\mid q_{mn}\neq 0\}\subset\mathbb{R}^{2},$
where the point (1,n)(-1,n) corresponds to the monomial $p_{0n}x_{2}^{n}$ and the point (m,1)(m,-1) corresponds to the monomial $q_{m0}x_{1}^{m}.$ Furthermore, $(0,0)$ is associated with the monomials $p_{10}x_{1}$ and $q_{01}x_{2}.$
Definition 2.5.
The Newton polygon associated to $f$ is given by the convex hull
$\bigcup_{(r,s)\in S}\{({r}^{*},{s}^{*})\in\mathbb{R}^{2}\mid r\leq{r}^{*},\,s\leq{s}^{*}\}.$
The Newton polygon associated to $f$ consists of the union $\gamma$ of the points/segments $\gamma_{k}$, enumerated in order from left to right. For each $1\leq k\leq n$, the segment $\gamma_{k}$ lies on the straight line $\alpha_{k}r+\beta_{k}s=d_{k}$ where $\alpha_{k}$ and $\beta_{k}$ are coprime integers and $d_{k}$ is a constant. Since $f$ has an isolated singularity at the origin, then at least one of the two points $(0,s)$ or (1,s)(-1,s) belongs to the support $S$ for some $s$ and, similarly, at least one of the two points (r,1)(r,-1) or $(r,0)$ lies in $S$, for some $r$. Thus a first segment $\gamma_{1}$ always exists in the Newton diagram ($\Leftrightarrow$ it is non-empty).
The Newton diagram identifies the dominant monomials near a singularity; the slopes of its compact edges determine the weight ratio $\beta/\alpha$, $\alpha\neq 0$, used to choose the exponents of a quasi-homogeneous blow-up.
Proposition 2.6.
[5, Section 3.3, adapted]. Under the previous conditions, the most efficient pair $(\alpha,\beta)$ for the quasi-homogeneous blow-up corresponds to the edge of the Newton polygon with minimal slope (in absolute value).
2.3. Poincaré compactification
Here, following Chapter 5 of [5], we present a concise overview of the Poincaré compactification to study how trajectories of a planar differential system behave “near infinity”. This technique will be used in Section 5 of the present paper.
To construct the phase portrait, one would ideally need to consider the complete real plane $\mathbb{R}^{2},$ which is often impractical. However, if the vector field is defined by polynomial functions, Poincaré compactification can be applied. This method allows the phase portrait to be represented within a finite region and also provides control over orbits that approach or come from infinity. Consider the polynomial vector field $f=(P,Q)$, where $d_{1}$ and $d_{2}$ are the algebraic degrees of $P$ and $Q$ respectively and $d={\max}\{d_{1},d_{2}\}.$
2.3.1. Preliminaries for the construction
First identify $\mathbb{R}^{2}$ with the plane $\Pi$ in $\mathbb{R}^{3}$ defined as $(y_{1},y_{2},y_{3})=(x_{1},x_{2},1).$ The sphere
$\mathbb{S}^{2}=\left\{(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}:\sum_{i=1}^{3}{y_{i}}^{2}=1\right\}$
is called the Poincaré sphere and is tangent to $\Pi$ at $(0,0,1).$ The sphere may be written as
$H^{+}\cup H^{-}\cup\mathbb{S}^{1}$
where
$H^{+}=\{(y_{1},y_{2},y_{3})\in\mathbb{S}^{2}:y_{3}>0\},\quad H^{-}=\{(y_{1},y_{2},y_{3})\in\mathbb{S}^{2}:y_{3}<0\}$
and the equator $\mathbb{S}^{1}$ given by:
$\mathbb{S}^{1}=\{(y_{1},y_{2},y_{3})\in\mathbb{S}^{2}:y_{3}=0\}.$
If $x=(x_{1},x_{2})\in\mathbb{R}^{2}$ define $\Delta(x)$ as $\sqrt{x_{1}^{2}+x_{2}^{2}+1}\neq 0.$
Definition 2.7.
The projection of $\mathbf{X}:\mathbb{R}^{2}\to\mathbb{S}^{2}$ is defined with the central projections
$f^{\pm}:\mathbb{R}^{2}\to\mathbb{S}^{2}$
where $f^{\pm}(x_{1},x_{2})$ is the intersection point of the line that passes through $(0,0)$ and $(x_{1},x_{2})\in\mathbb{R}^{2}$, with $H^{\pm}$. More precisely:
$f^{+}(x_{1},x_{2})=\left(\frac{x_{1}}{\Delta(x)},\frac{x_{2}}{\Delta(x)},\frac{1}{\Delta(x)}\right).$
Analogously, we define $f^{-}$ (substituting $H^{+}$ by $H^{-}$) as:
$f^{-}(x_{1},x_{2})=\left(-\frac{x_{1}}{\Delta(x)},-\frac{x_{2}}{\Delta(x)},-\frac{1}{\Delta(x)}\right).$
For $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, the induced vector field on $H^{\pm}$ is defined, respectively, by:
$\overline{\mathbf{X}}(y)$ $=Df^{\pm}(x)\mathbf{X}(x),\quad\text{for}\quad y=f^{\pm}(x).$
Remark 2.8.
The vector field $\overline{\mathbf{X}}$ is tangent to the sphere $\mathbb{S}^{2}$ at every point $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$.
The infinite points of $\mathbb{R}^{2}$ (each direction is associated with two points) are in bijective associated to the points of the equator of $\mathbb{S}^{2}.$ In the next subsection, we are going to generalize the induced vector field $\overline{\mathbf{X}}:\mathbb{S}^{2}\backslash\mathbb{S}^{1}\rightarrow\mathbb{S}^{2}.$
Definition 2.9.
The extension of $\mathbf{X}$ to $\mathbb{S}^{2}$ is referred to as its Poincaré compactification, denoted by $p(\mathbf{X}).$
2.3.2. About the construction
We use smooth charts to make calculations. For $y=(y_{1},y_{2},y_{3})\in\mathbb{S}^{2},$ we utilize the six local charts:
$U_{l}=\{y\in\mathbb{S}^{2}:0<y_{l}\},\quad V_{l}=\{y\in\mathbb{S}^{2}:0>y_{l}\},\quad l=1,2,3.$
The associated local maps $\Phi_{l}:U_{l}\to\mathbb{R}^{2}$ and $\Psi_{l}:V_{l}\to\mathbb{R}^{2}$ are described as
$\Phi_{l}(y)=-\Psi_{l}(y)=\left(\frac{y_{m}}{y_{l}},\frac{y_{n}}{y_{l}}\right),\quad\text{for}\quad m<n\quad\text{where}\quad m,n\neq l.$
Consider $z=(u,v)$ as the value of $\Phi_{l}(y)$ or $\Psi_{l}(y)$ for any $l\in\{1,2,3\}$, such that $(u,v)$ plays different roles depending on the local chart we are considering. In any chart, points on $\mathbb{S}^{1}$ correspond to those with $v=0$.
2.3.3. Explicit expressions
In the local chart $(U_{1},\Phi_{1}),$ $p(\mathbf{X})$ is expressed as
$\dot{u}=\left[Q\left(\frac{1}{v},\frac{u}{v}\right)-uP\left(\frac{1}{v},\frac{u}{v}\right)\right]v^{d},\hskip 14.22636pt\dot{v}=-P\left(\frac{1}{v},\frac{u}{v}\right)v^{d+1},$ (2.3)
the expression for $(U_{2},\Phi_{2})$ is:
$\dot{u}=\left[P\left(\frac{u}{v},\frac{1}{v}\right)-uQ\left(\frac{u}{v},\frac{1}{v}\right)\right]v^{d},\hskip 14.22636pt\dot{v}=-Q\left(\frac{u}{v},\frac{1}{v}\right)v^{d+1},$ (2.4)
and the expression for $(U_{3},\Phi_{3})$ is:
$\dot{u}=P(u,v),\hskip 14.22636pt\dot{v}=Q(u,v).$ (2.5)
Remark 2.10.
For $i=1,2,3$, the representations for the vector field $p(\mathbf{X})$ in the local chart $(V_{i},\Psi_{i})$ are identical to those in $(U_{i},\Phi_{i}),$ up to a multiplication by $({-1})^{d-1}$.
To investigate $\mathbf{X}$ throughout the complete plane $\mathbb{R}^{2}$, as well as its behaviour “near infinity”, it is enough to examine it on $H^{+}\cup\mathbb{S}^{1}$, a region known as the Poincaré disc. All calculations can be performed in the three charts $(U_{i},\Phi_{i})$ for $i=1,2,3$ whose expressions are given by the formulas (2.3), (2.4) and (2.5).
Definition 2.11.
We refer to singular points of $\mathbf{X}$ as finite when the corresponding points of $p(\mathbf{X})$ are in $\mathbb{S}^{2}\setminus\mathbb{S}^{1},$ and as infinite when they belong to $\mathbb{S}^{1}$.
2.3.4. Useful consequences
Following Chapter 5 of [5], we list some useful consequences of the theory described above.
  1. (1)
    Whenever $y\in\mathbb{S}^{1}$ is an infinite singular point, so is y-y.
  2. (2)
    The local behaviour near y-y is that near $y$ scaled by $(-1)^{d-1}$.
  3. (3)
    Because infinite singular points arise in pairs, it is sufficient to examine just half of them. The rest can be inferred via the vector field’s degree.
  4. (4)
    The integral curves on $\mathbb{S}^{2}$ are symmetric aroud the origin, so it suffices to depict the flow of $p(\mathbf{X})$ in the closed northern hemisphere, the so called Poincaré disc.
2.3.5. Infinite singular points
The theory described here will be useful in Section 5 of the present paper. Our goal is to analyse the local phase portrait near infinite singular points. To this end, we select a singular point $(u,0)$ and begin with examining the linear part of the vector field $p(\mathbf{X})$. Denote by $P_{i}$ and $Q_{i}$ the homogeneous polynomials of degree $i\in\mathbb{N}_{0}$ for i=0,1,,di=0,1,...,d such that
$\left\{\begin{array}[]{l}P=P_{0}+P_{1}+...+P_{d},\\ \\ Q=Q_{0}+Q_{1}+...+Q_{d}.\end{array}\right$.
Then $(u,0)\in(U_{1}\cup V_{1})\cap\mathbb{S}^{1}$ is an infinite singular point of $p(\mathbf{X})$ if and only if
$F(u)\equiv-uP_{d}(1,u)+Q_{d}(1,u)=0.$
In the same way $(u,0)\in(U_{2}\cup V_{2})\cap\mathbb{S}^{1}$ is an infinite singular point for $p(\mathbf{X})$ if and only if
$G(u)\equiv-uQ_{d}(u,1)+P_{d}(u,1)=0.$
The Jacobian of the vector field $p(\mathbf{X})$ at the point $(u,0)$ is
$\left(\begin{array}[]{cc}F^{\prime}(u)&Q_{d-1}(1,u)-uP_{d-1}(1,u)\\ \\ 0&-P_{d}(1,u)\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{cc}G^{\prime}(u)&P_{d-1}(u,1)-uQ_{d-1}(u,1)\\ \\ 0&-Q_{d}(u,1)\end{array}\right)$
if $(u,0)$ lies in $U_{1}\cup V_{1}$ or $U_{2}\cup V_{2}$, respectively.
Our discussion is concentrated on isolated singularities in the equator. According to [5], only nodes and saddles can arise as hyperbolic singularities at infinity. All semi-hyperbolic singular points may also emerge at infinity. In the case that one of these singular points at infinity is a (topological) saddle, the line $v=0,$ corresponding to the equator of $\mathbb{S}^{2}$, serves as a stable, unstable, or center manifold. This property also applies to semi-hyperbolic points of saddle-node type, whose hyperbolic sectors may be divided into two different forms depending on the Jacobian matrix in charts $U_{1}$ or $U_{2}.$ These matrices may be either
$\left(\begin{array}[]{cc}0&\star\\ \\ 0&a\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{cc}a&\star\\ \\ 0&0\end{array}\right)$
where $a\neq 0$ and $\star\in\mathbb{R}$. Nilpotent equilibria exhibit behaviour at infinity that is considerably more complex than those of hyperbolic equilibria. Blow-up techniques of Subsection 2.2 are needed to study them.
3. Finite equilibria and bifurcation analysis
Our object of study is the analysis of the family of differential equations (1.5), which can be written as:
$\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}(x+by-a)\end{array}\right$. (3.1)
[…]
$Df_{(a,b,c)}(x,y)=\left(\begin{array}[]{cc}c-cx^{2}&c\\ \\ -1/c&-b/c\end{array}\right).$ (3.5)
It is immediate to deduce:
Lemma 3.1.
The divergence of (3.1) is given by $c-cx^{2}-b/c,$ which is strictly negative for $b>c^{2}.$
By the Bendixson criterion [5, Theorem 7.10], the negativity of the previous result ensures that the system (3.1) cannot have periodic orbits and, consequently, no limit cycles in the open set defined by the inequality $b>c^{2}.$ From now on, we divide the analysis into Cases A, B and C (see Subsection 1.4), depending on the parameters $a,b$ and $c$ of (3.1).
3.1. Case A ($a=0$)
Consider the invertible linear map in $\mathbb{R}^{2}$ defined by $\kappa(x,y)=(-x,-y)$, whose action on $\mathbb{R}^{2}$ is isomorphic to that of $\mathbb{Z}_{2}$ (rotation of $\pi$ around the origin). Since
$f_{(0,b,c)}\circ\kappa=\kappa\circ f_{(0,b,c)},$
we may say that:
Lemma 3.2.
The vector field $f_{(0,b,c)}$ is $\mathbb{Z}_{2}(\kappa)$–equivariant.
[…]
$E_{2}=\kappa(E_{3})=\left(-\sqrt{\dfrac{3(b-1)}{b}},\dfrac{1}{b}\sqrt{\dfrac{3(b-1)}{b}}\right)$ (3.6)
and $E_{3}=\kappa(E_{2})=\left(\sqrt{\dfrac{3(b-1)}{b}},-\dfrac{1}{b}\sqrt{\dfrac{3(b-1)}{b}}\right).$
From (3.5), we deduce that the trace and determinant operators of $Df_{(0,b,c)},$ evaluated at $E_{i},$ $i=1,2,3,$ are given and denoted, respectively, by:
$\mbox{Tr}(E_{1})$ $=c-\dfrac{b}{c},\qquad\qquad\qquad\qquad\mbox{Det}(E_{1})=1-b,$
$\mbox{Tr}(E_{2,3})$ $=-\dfrac{2c^{2}b-3c^{2}+b^{2}}{cb},\qquad\mbox{Det}(E_{2,3})=2(b-1).$
It can be easily shown that $E_{1}$ is a hyperbolic saddle for $b>1$ (since $\mbox{Det}(E_{1})<0$). Using (3.5), the jacobian matrix of the vector field $f_{(0,1,\pm 1)}$ at $E_{1}\equiv E_{2}\equiv E_{3}$ is given by:
$\left(\begin{array}[]{cc}\pm 1&\pm 1\\ \\ \mp 1&\mp 1\end{array}\right).$ (3.10)
It is straightforward to see that the matrix (3.10) is non-hyperbolic and has a double zero eigenvalue. Our main result relies on the existence of a Double-zero bifurcation (DZ) with $\mathbb{Z}_{2}(\kappa)$–symmetry of $f_{(0,1,\pm 1)}$ at the singularity $E_{1}$.
Remark 3.3.
Consider the following linear change of coordinates, $\Psi,$ and its inverse:
$\Psi(x,y)=(x+y,-x)\quad\text{and}\quad\Psi^{-1}(x,y)=(-y,x+y).$
One may easily check that:
$\Psi^{-1}\circ f_{(0,1,\pm 1)}\circ\Psi(x,y)$ $=$ $\Psi^{-1}\circ f_{(0,1,\pm 1)}(x+y,-x)$
$=$ $\Psi^{-1}\left(-\frac{(x\pm y)^{3}}{3}+y,-y\right)$
$=$ $\left(y,-\frac{1}{3}(x\pm y)^{3}\right),$
so that the equations in Jordan form are represented by
$\left\{\begin{array}[]{l}\dot{x}=y\\ \\ \dot{y}=-\frac{1}{3}(x\pm y)^{3},\end{array}\right$. (3.11)
which contains part of the truncated form of degree 3 of the versal deformation of a Double-zero bifurcation with $\mathbb{Z}_{2}$–symmetry [27, pp. 400, Case c=1c=-1]. The next result deals with the description of all bifurcation curves passing through the bifurcation point $(b,c)=(1,\pm 1)$ and that characterises a DZ bifurcation of $f_{(0,b,c)}$ with $\mathbb{Z}_{2}(\kappa)$–symmetry.
We focus the proof on the analysis around the bifurcation parameter $(b,c)=\left(1,1\right)$; the analysis close to $(b,c)=\left(1,-1\right)$ has a similar treatment (reversing time).
Theorem 3.4.
Regarding to the vector field $f_{(0,b,c)},$ the equilibrium $E_{1}\equiv E_{2}\equiv E_{3}$ undergoes a DZ bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1).$ The global representations of the transition/bifurcation curves in the parameters space $(b,c)\in\mathbb{R}\times\mathbb{R}^{+}$ are as follows (schematic lines for $a=0$ have been plotted in Figures 1 and 3):
  1. (1)
    The equilibrium $E_{1}$ undergoes a supercritical (with respect to $b$) Pitchfork bifurcation along the line:
    $T_{P}=\{(b,c):b=1\}$
    giving rise to the equilibria $E_{2}$ and $E_{3}$ defined in (3.6) for $b>1$.
  2. (2)
    The equilibrium $E_{1}$ undergoes a Belyakov transition along the union of parabola:
    $T^{1}_{F}=\{(b,c):b=-c^{2}\pm 2c\}.$
  3. (3)
    The equilibrium $E_{1}$ undergoes a supercritical (with respect to $b$) Hopf bifurcation along part of the parabola:
    $T_{H}^{1}=\left\{(b,c):b=c^{2},\,|c|<1\right\}.$
  4. (4)
    The equilibria $E_{2}$ and $E_{3}$ undergo a Belyakov transition along the algebraic curves:
    $T^{2,3}_{F}=\left\{(b,c):\left(b^{2}+2cb-2c^{2}b+3c^{2}\right)\left(b^{2}-2cb-2c^{2}b+3c^{2}\right)=0\right\}.$
  5. (5)
    The equilibria $E_{2}$ and $E_{3}$ undergo a subcritical (with respect to $b$) Hopf bifurcation along the algebraic curves:
    $T^{2,3}_{H}=\left\{(b,c):b=c\left(-c\pm\sqrt{c^{2}+3}\right)\mbox{ and }b>1\right\}.$
  6. (6)
    There is a double homoclinic cycle to $E_{1}$ along the approximated curve:
    $DH:\quad 0=\frac{7b^{2}+10bc^{2}-17c^{2}}{15c^{3}}\sqrt{\frac{-7b^{2}+5bc^{2}+2c^{2}}{5b}},\quad b\in(0,c).$
Remark 3.5.
For every pair $(b,c)\in\mathbb{R}\times\mathbb{R}^{+}$, the origin is always an equilibrium point. When $b<0$, two additional nontrivial equilibria appear; they move away from the origin and diverge to infinity as $b\to 0^{-}$. For $0<b<1$, the origin is the only equilibrium. Thus, although the number of equilibria changes as the line $b=0$ is crossed, this change results from equilibria escaping to (or emerging from) infinity rather than from a local bifurcation at the origin.
Remark 3.6.
According to [21, Theorem 3.5.1] one knows that there is a Bautin bifurcation along the curve parametrized by $c$ defined by
$\left\{(a,b,c)\in\mathbb{R}^{2}\times\mathbb{R}^{+}:\left(\pm\frac{4}{3}\left(c\sqrt{c^{2}-1}-c^{2}+1\right)\sqrt[4]{1-\frac{1}{c^{2}}},c^{2}\pm c\sqrt{c^{2}-1},c\right),c\geq 1\right\}$ (3.12)
In particular, in the three-dimensional bifurcation space $(a,b,c)$, this gives rise to a saddle-node bifurcation surface associated with two non-hyperbolic cycles, generating the line SNL defined by $b=c^{2}\pm c\sqrt{c^{2}-1}$ and $a=0$ of Figure 6. The formula of [21] is explicitly written for $c\geq 1$, however due to the invariance under the reversible transformation $(t,c)\to(-t,-c)$, it also holds for $c\leq-1$.
The curve SNL of Figure 6 does not make part of a generic unfolding of a DZ bifurcation with symmetry; this is why it is represented as a dashed line.
When $a=0$, the points $\textbf{Q}_{17}$ and $\textbf{Q}_{18}$ of [21, pp. 180] coincide with the point $(b,c)=(1,1)$. This coincidence yields a codimension-three DZ–bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1)$, whose analysis is beyond the scope of the present paper. A complete understanding of this bifurcation remains an open problem and is briefly discussed in Section 6.
Figure 1. Illustration of Theorem 3.4. Left: The bifurcation curves associated to (3.1) for Case A. Linear stability of $E_{1}$ for $f_{(0,b,c)}$. Right: Linear stability of $E_{2}$ and $E_{3}$ for $f_{(0,b,c)}$. The notation $T_{P}$, $T_{H}^{1}$, $T_{F}^{1}$, $T_{P}$ $T_{H}^{2,3}$, $T_{F}^{2,3}$ follows from Theorem 3.4 and sd, uf, un, sf, and sn refer to saddle, unstable focus, unstable node, stable focus and stable node.
Remark 3.7.
Theorem 3.4 generalizes the analysis of [8] and [21] by locating a codimension 2 bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1)$ for the vector field $f_{(0,b,c)}.$ For the sake of completeness, we present local numerical phase portraits on Table 1 of Appendix. All of them agree well with the theory.
Proof of Theorem 3.4.
The first statement of the result comes from the normal form (3.11). The eigenvalues of the Jacobian matrix (3.5) at $E_{1},$ $E_{2}$ and $E_{3},$ when the equilibria exist, are:
$E_{1}:\quad$ $\lambda_{\pm}^{1}=\dfrac{c^{2}-b\pm\sqrt{(c^{2}+b)^{2}-4c^{2}}}{2c},$
$E_{2},E_{3}:\quad$ $\lambda_{\pm}^{2,3}=\frac{3{c}^{2}-{b}^{2}-2{c}^{2}b\pm\sqrt{4{c}^{4}{b}^{2}-12{c}^{4}b-4{c}^{2}{b}^{3}+9{c}^{4}+2{c}^{2}{b}^{2}+{b}^{4}}}{2cb}.$
Based on linear analysis, the type of equilibria for (3.1) ($a=0$) and for $b,c\in[-4,4]\times[-2,2]$ in different regions defined by (3.1) are provided in Figures 1 and 3.
  1. (1)
    We prove the existence of a Pitchfork bifurcation for $b=1$ using the normal form truncated at order 3 [27, Section 20.1E]. Let $b=1+\mu$, where $\mu\in\mathbb{R}$ is an additional bifurcation parameter. For $\mu=0$, the Jacobian matrix (3.5) associated with $E_{1}$ admits eigenvalues: 0 and $\frac{c^{2}-1}{c}$, $c\neq 0,1$. Define the diffeomorphism $T:\mathbb{R}^{2}\to\mathbb{R}^{2}$ as
    $\mathbf{x}\equiv(x,y)\mapsto\left(\frac{c^{2}y-x}{c^{2}-1},-\frac{y-x}{c^{2}-1}\right).$
    In the new coordinates $\mathbf{x}=(x,y),$ where $\mathbf{0}=(0,0)$, let
    $\bar{f}_{(0,b,c)}(\mathbf{x})\equiv f_{(0,b,c)}(\mathbf{x})-Df_{(0,b,c)}(\mathbf{0})\mathbf{x}.$
    Therefore, we have
    $\dot{\mathbf{x}}=$ $T^{-1}Df_{(0,b,c)}(\mathbf{0})T\mathbf{x}+T^{-1}\bar{f}_{(0,b,c)}T\mathbf{x}$ (3.13)
    $=$ $\begin{pmatrix}1&c^{2}\\ 1&1\end{pmatrix}\begin{pmatrix}c&c\\ -\frac{1}{c}&-\frac{1}{c}\end{pmatrix}\begin{pmatrix}\frac{1}{1-c^{2}}&\frac{c^{2}}{c^{2}-1}\\ \frac{1}{c^{2}-1}&\frac{1}{1-c^{2}}\end{pmatrix}$+
    $\begin{pmatrix}-cy\mu-\frac{1}{3}cx^{3}\\ \frac{c^{2}-1}{c}x+\frac{c^{2}-\mu-1}{c}y-\frac{c}{3}x^{3}\end{pmatrix}\Big|_{x\rightarrow\frac{c^{2}y-x}{c^{2}-1},y\rightarrow\frac{x-y}{c^{2}-1}}$
    from where we conclude that (in the new coordinates above):
    $\left\{\begin{array}[]{l}\dot{x}=-\frac{c}{3(c^{2}-1)^{3}}\Big(3c^{2}x^{2}y-x^{3}-3c^{4}xy^{2}+c^{6}y^{3}+3(c^{2}-1)^{2}\mu x-3(c^{2}-1)^{2}\mu y\Big),\\ \\ \dot{y}=-\frac{1}{c(c^{2}-1)^{3}}\Big(3c^{4}x^{2}y-c^{2}x^{3}-3c^{6}xy^{2}+c^{8}y^{3}+3(c^{2}-1)^{2}\mu x-3(c^{2}-1)^{2}\mu y-3(c^{2}-1)^{4}y\Big).\end{array}\right$.
    We aim to find a center manifold in the form of
    $h(x,\mu):={\alpha_{0}}\mu^{2}+\alpha_{1}x^{2}+\alpha_{2}x\mu+\dots.$
    Since the center manifold must satisfy the equation (cf. [27, Equation 3.2.7])
    $\dot{x}\frac{\partial h}{\partial x}-\dot{y}\big|_{(x,h(x,\mu))}=0,$
    we obtain $\alpha_{0}=\alpha_{1}=0,$ and $\alpha_{2}=\frac{1}{\left(c^{2}-1\right)^{2}}.$ Then, the dynamics of (3.13) restricted to the center manifold $y=\frac{\mu x}{(c^{2}-1)^{2}}$ is given by:
    $\dot{x}=-\frac{c}{c^{2}-1}\mu x+\frac{c}{3\left(c^{2}-1\right)^{3}}x^{3}$ (3.14)
    meaning that there is a Pitchfork bifurcation at $\mu=0$. If $c>1$, then the bifurcation is supercritical with respect to $\mu$ and also with respect to $b$ in the original equation, as depicted in Figure 2. The analysis for the case $c>1$ is different from the case $0<c<1$, although the argument runs along the same lines.
    Figure 2. Illustration of the Pitchfork bifurcation of (3.14) for $c>1$ and close to $\mu=0(\Leftrightarrow b=1)$. (a) $\mu<0$, (b) $\mu=0$ and (c) $\mu>0$.
  2. (2)
    We should prove that the eigenvalues of the Jacobian matrix of $f_{(0,b,c)}$ at $E_{1}$ change from real to complex (non-real). The proof of this item follows by observing that the eigenvalues associated with the Jacobian matrix (3.5) at $E_{1}$ are complex (non-real) for $b>-c^{2}+2c$ or $b<-c^{2}-2c,$ and real otherwise. The sign of their real part does not change.
  3. (3)
    Let $b=c^{2}$ and $|c|<1$. The Jacobian matrix of $f_{(0,b,c)}$ at $E_{1}$ has a pair of purely imaginary eigenvalues $i\sqrt{1-c^{2}}$ with 1<c<1.-1<c<1. Let $b=c^{2}+\mu$ where $\mu\in\mathbb{R}$ is a bifurcation parameter. The eigenvalues associated with the Jacobian matrix at $E_{1}$ for $\mu\neq 0$
    $\begin{pmatrix}c&c\\ -\frac{1}{c}&-c-\frac{1}{c}\mu\end{pmatrix}$,
    are given by
    $\lambda(\mu)=-\frac{1}{2c}\mu\pm\frac{1}{2c}\sqrt{(\mu+2c^{2}+2c)(\mu+2c^{2}-2c)}.$ (3.15)
    For $\mu$ close to zero and $|c|<1$, there exists a pair of complex conjugate eigenvalues. Furthermore, since
    $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}=-\frac{1}{2c}\neq 0,$
    this means that there is a non-degenerate Hopf bifurcation at $\mu=0$ ([17, pp. 299–300]). By using the linear transformation
    $\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}&\dfrac{c}{2\sqrt{1-c^{2}}}\\ 0&-\dfrac{1}{2c\sqrt{1-c^{2}}}\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}$,
    the system (3.1) turns into its Jordan form
    $\left\{\begin{array}[]{l}f^{1}:\,\dot{u}=-\frac{1}{12}c{u}^{3}-\frac{1}{4}\frac{{c}^{2}}{\sqrt{1-{c}^{2}}}v{u}^{2}-\frac{1}{4}{\frac{{c}^{3}}{1-{c}^{2}}}{v}^{2}u-\frac{1}{12}\frac{{c}^{4}}{\left(1-{c}^{2}\right)^{3/2}}{v}^{3}-\sqrt{1-{c}^{2}}v,\\ \\ f^{2}:\,\dot{v}=\sqrt{1-{c}^{2}}u.\end{array}\right$.
    The first Lyapunov coefficient depends on $c$ and is given by:
    $\ell(c)=\dfrac{1}{16\omega}\left(R_{1}+\omega R_{2}\right),$
    where
    $\omega:=$ $\sqrt{1-{c}^{2}}\neq 0$
    $R_{1}:=$ $f^{1}_{uv}\left(f^{1}_{uu}+f^{1}_{vv}\right)-f^{2}_{uv}\left(f^{2}_{uu}+f^{2}_{vv}\right)-f^{1}_{uu}f^{2}_{uu}+f^{1}_{vv}f^{2}_{vv}\quad\text{and}$
    $R_{2}:=$ $f^{1}_{uuu}+f^{1}_{uvv}+f^{2}_{uuv}+f^{2}_{vvv}.$
    Using Maple, we get
    $\ell(c)=\frac{c}{32(c^{2}-1)}<0$
    and we may conclude that the limit cycle is stable for $0<c<1$ and the bifurcation is supercritical – [27, Remark 1, pp. 383]. In this case, $E_{1}$ is an asymptotically stable equilibrium for $\mu>0$ ($\Leftrightarrow b>c^{2}$) and unstable for $\mu<0$ ($\Leftrightarrow b<c^{2}$), with an asymptotically stable periodic solution for $\mu<0$ ($\Leftrightarrow b<c^{2}$).
  4. (4)
    The proof is similar as item (2) by taking into account that
    $4{c}^{4}{b}^{2}-12{c}^{4}b-4{c}^{2}{b}^{3}+9{c}^{4}+2{c}^{2}{b}^{2}+{b}^{4}$
    $=\left(b^{2}+2cb-2c^{2}b+3c^{2}\right)\left(b^{2}-2cb-2c^{2}b+3c^{2}\right).$
  5. (5)
    Using linear analysis, one knows that Hopf bifurcation exists only if $\mbox{Tr}(E_{2})=\mbox{Tr}(E_{3})=0$ and $\mbox{Det}(E_{2}),\mbox{Det}(E_{3})>0.$ Indeed, we have:
    $\mbox{Tr}(E_{2})=\mbox{Tr}(E_{3})=0\mbox{ and }b>1$
    $\Leftrightarrow$ $3{c}^{2}-{b}^{2}-2{c}^{2}b=0\mbox{ and }b>1$
    $\Leftrightarrow$ $b=c\left(-c\pm\sqrt{c^{2}+3}\right)\mbox{ and }b>1.$
    Figure 3. The bifurcation curves associated to (3.1) for Case A. The local portraits corresponding to Regions 1–28 have been plotted in Table 1 of Appendix.
    The inequality $b>1$ serves to ensure that the equilibria $E_{2}$ and $E_{3}$ exist.
    Let $b=c\left(-c\pm\sqrt{c^{2}+3}\right).$ The Jacobian matrix at $E_{2,3}$ has a pair of purely imaginary eigenvalues $\pm i\sqrt{2c\sqrt{c^{2}+3}-2-2c^{2}}$ with $c>1.$ Let $b=-c^{2}\pm c\sqrt{c^{2}+3}+\mu$ where $\mu\in\mathbb{R}$ denotes the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ associated with the Jacobian matrix at $E_{2,3}$
    $\begin{pmatrix}-\frac{c\left(2{c}^{2}+2c\sqrt{{c}^{2}+3}\pm 2\mu\mp 3\right)}{\mp{c}^{2}+c\sqrt{{c}^{2}-3}\pm\mu}&c\\ -\frac{1}{c}&{\frac{{c}^{2}\mp c\sqrt{{c}^{2}+3}-\mu}{c}}\end{pmatrix}$,
    are given by (computations performed in Maple):
    $\lambda(\mu)=\dfrac{-2c\sqrt{c^{2}+3}\mu-\mu^{2}\pm\sqrt{\sigma}}{2c(\mu-c^{2}+c\sqrt{c^{2}+3})}$
    where
    $\sigma=\mu^{4}+4c(\sqrt{c^{2}+3}-2c)\mu^{3}-4c^{2}(6c\sqrt{c^{2}+3}-7c^{2}-5)\mu^{2}$
    $+8c^{3}(2\sqrt{c^{2}+3}(3c^{2}+1)-6c^{3}-11c)\mu-8c^{4}(c\sqrt{c^{2}+3}\left(4c^{2}+5\right)-11c^{2}-4c^{4}-3).$
    There is a pair of complex conjugate eigenvalues when $\sigma<0$. In addition, we have:
    $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}=-\frac{\sqrt{c^{2}+3}}{c(\sqrt{c^{2}+3}\mp c)}.$
    The first Lyapunov coefficient $\ell$ associated to $f_{(0,b,c)}$ at $E_{2,3}$ depends on $c$ and is given explicitly by (computations in Maple):
    $\ell(c)={\frac{c\left(\left(4c^{2}+3\right)\left(\sqrt{c^{2}+3}-c\right)-3c\right)\left(\sqrt{c^{2}+3}c\left(21+44c^{2}+16c^{4}\right)-c^{2}\left(69+68c^{2}+16c^{4}\right)-9\right)}{32\left(\sqrt{c^{2}+3}c-c^{2}-1\right)^{2}\left(\sqrt{c^{2}+3}-c\right)^{7}}},$
    Since $\lim_{c\to 1^{+}}\ell(c)=\lim_{c\to+\infty}\ell(c)=+\infty$ and $\frac{\partial\ell(c)}{\partial c}=0\quad\text{at}\quad c^{\star}\approx 1.315$, then $\ell$ remains positive in the neighbourhood of $c^{\star}$. Therefore, the corresponding limit cycles for $c>1$ are unstable and the bifurcation is subcritical.
  6. (6)
    This item follows from [8, pp. 201] and [21, Formula (3.2.29)].
Remark 3.8.
Numerically, we may observe that the non-degenerate Hopf bifurcation of items (3) and (5) of Theorem 3.4 are, respectively, supercritical and subcritical. The bifurcation $T_{H}^{1}$ yields a stable periodic solution, say $C^{s}$, and $T_{H}^{2,3}$ generates two unstable and $\kappa$–symmetric periodic solutions – confirm on Table 1 of Appendix.
3.2. Case B ($b=0$)
Model (3.1) for $b=0$ has only one equilibrium given explicitly by:
$E_{1}=\,\left(a,-a+\frac{1}{3}a^{3}\right),$
and the linear part of $f_{(a,0,c)}$ at $E_{1}$ has eigenvalues
$\frac{1}{2}c(1-a^{2})\pm\frac{1}{2}\sqrt{(ca^{2}-c+2)(ca^{2}-c-2)}.$ (3.16)
Then we may conclude that:
Theorem 3.9.
[…]
  • I.
    The equilibrium $E_{1}$ undergoes a Belyakov transition along the hyperbola:
    $T^{b0}_{F}=\left\{(a,c):c=\pm\frac{2}{a^{2}-1}\right\}.$
    Figure 4. The classification of the equilibrium point $E_{1}$ and the corresponding bifurcation curves for (3.1) in Case B. In this figure, un, sn, uf and sf refer to unstable (stable) node and unstable (stable) focus, respectively. The local phase portraits associated with Regions 1–8 have been plotted in Table 2.
  • II.
    The equilibrium $E_{1}$ undergoes a Hopf bifurcation along the lines:
    $T^{b0}_{H}=\left\{(a,c):a=\pm 1\right\}.$
    The Hopf bifurcation is supercritical(resp. subcritical) if $c>0$ (resp. $c<0$).
Schematic bifurcation lines for $b=0$ have been plotted in Figure 4. The local phase portraits associated with Regions 1–8 have been plotted in Table 2
Proof.
We are going to concentrate the proof in the case $b=0$ and $c>0$. Again, the analysis of the case $b=0$ and $c<0$ has a similar treatment (reversing time).
  • I.
    The first claim is a direct consequence of (3.16).
  • II.
    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues $\pm i.$ Let $a=\pm 1+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ corresponding to the Jacobian matrix (3.5) at $E_{1}$ are (computations performed with Maple):
    $\lambda(\mu)=\mp c\mu-\frac{1}{2}c{\mu}^{2}\pm\frac{1}{2}\sqrt{4{c}^{2}{\mu}^{2}\pm 4{c}^{2}{\mu}^{3}+{c}^{2}{\mu}^{4}-4}.$
    For $\mu$ close to zero, we have $4{c}^{2}{\mu}^{2}+4{c}^{2}{\mu}^{3}+{c}^{2}{\mu}^{4}-4<0$, which means that $\lambda(\mu)$ is complex non-real. Furthermore
    $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}\big|_{\mu=0}=\mp c\neq 0.$
    The first Lyapunov coefficient depends on $c$ and is given by
    $\ell(c)=-\frac{1}{32}c$
    which has opposite sign to that of $c\neq 0$. Hence the bifurcation is subcritical for $c<0$ and supercritical for $c>0$.
    For the case a=1a=-1, $E_{1}$ is an asymptotically stable equilibrium for $\mu<0$ and unstable focus for $\mu>0$ with an asymptotically stable periodic solution emerging for $\mu>0$. For the case $a=1$, $E_{1}$ is an asymptotically stable equilibrium for $\mu>0$ and unstable focus for $\mu<0$ with an asymptotically stable periodic solution emerging $\mu>0$. This finishes the proof.
3.3. Case C ($a>0,$ $0<b<1$)
If $(x^{\star},y^{\star})$ is an equilibrium of (3.1) under the conditions $a>0,$ $0<b<1,$ then:
$x^{\star}-a+by^{\star}=0\Leftrightarrow y^{\star}=\frac{a-x^{\star}}{b}.$
By substituting $y^{\star}=\frac{a-x^{\star}}{b}$ into the first equation of (3.1), then $x^{\star}$ is a root of $H(x)=0,$ where
$H(x)=-\frac{1}{3}cx^{3}+c\left(1-\frac{1}{b}\right)x+\frac{ac}{b}.$
The discriminant associated with the polynomial $H(x)$ is given by
$\Delta=\frac{4}{3}c^{4}\left(1-\frac{1}{b}\right)^{3}-3\frac{c^{4}a^{2}}{b^{2}}.$
For $0<b<1$, the discriminant satisfies $\Delta<0$. According to Cardano’s formula, this implies that the polynomial $H(x)$ has exactly one real root, denoted by $x^{\star}.$ Then the equilibrium for system (3.1) is explicitly given by
$\left(x^{\star},\frac{a-x^{\star}}{b}\right)=(x^{\star},y^{\star})=:E_{1}.$
The characteristic equation of the Jacobian matrix evaluated at $(x^{*},y^{*})$ is
$\lambda^{2}+\frac{c^{2}{x^{*}}^{2}-c^{2}+b}{c}\lambda+b{x^{*}}^{2}-b+1=0.$
We may then conclude that:
Proposition 3.10.
The following assertions hold for the vector field $f_{(a,b,c)}$ of (3.1), under the conditions $a>0,$ and $0<b<1$:
  1. (1)
    The equilibrium point $E_{1}$ is stable if $\frac{c^{2}{x^{*}}^{2}-c^{2}+b}{c}>0.$ Conversely, it is unstable if $\frac{c^{2}{x^{*}}^{2}-c^{2}+b}{c}<0.$
  2. (2)
    A non-degenerate Hopf bifurcation occurs at $c^{2}=\frac{b}{1-{x^{*}}^{2}}.$ The numerical simulations indicate the resulting limit cycle is attracting when $c>\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and repelling for $c<-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
Note that $x^{\star}\neq\pm 1$ (if $x^{\star}=\pm 1$ then $b=0$, which is a contradiction).
As before, we use the notation $\mbox{Det}(E_{1})$ and $\mbox{Tr}(E_{1})$ for the determinant and trace operators of $Df_{(a,b,c)},$ evaluated at $E_{1}$.
Proof.
Under the condition $0<b<1,$ one knows that $\det(E_{1})>0.$
  1. (1)
    The result follows by noticing that $\mbox{Tr}(E_{1})=-\frac{c^{2}{x^{*}}^{2}-c^{2}+b}{c}.$ If $\mbox{Tr}(E_{1})<0$, then the equilibrium $E_{1}$ is stable. Conversely, it is unstable if $\mbox{Tr}(E_{1})>0$.
  2. (2)
    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues $\pm i\sqrt{b{x^{*}}^{2}-b+1}.$ Let $b=c^{2}\left(1-{x^{*}}^{2}\right)+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ corresponding to the Jacobian matrix at $E_{1},$
    $\begin{pmatrix}c(1-x^{*})&c\\ -\frac{1}{c}&c{x^{*}}^{2}-c-\frac{\mu}{c}\end{pmatrix}$,
    are
    $\lambda(\mu)=\frac{-\mu\pm\sqrt{\left(2c^{2}{x^{*}}^{2}-2c^{2}-\mu-2c\right)\left(2c^{2}{x^{*}}^{2}-2c^{2}-\mu+2c\right)}}{2c}.$
    If
    $\min_{c\geq 0}{c\left(\pm 1-c+c{x^{*}}^{2}\right)}<\mu<\max_{c\geq 0}{c\left(\pm 1-c+c{x^{*}}^{2}\right)},$
    then the expression under the square root is negative. Hence there exists a pair of complex non real conjugate eigenvalues. The result follows by observing that (computations performed with Maple):
    $\frac{d{\rm Re}\,{\lambda(\mu)}}{d\mu}=-\frac{1}{2c}\neq 0.$
Remark 3.11.
The statement of Proposition 3.10 says nothing about the criticality of the Hopf Bifurcation. Analytically, it is very difficult to compute the first Lyapunov coefficient of (3.1) at an equilibrium whose explicit expression we do not know. However, numerics suggest that the Hopf bifurcation is supercritical for $c=\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and subcritical for $c=-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
4. Analysis of the singular case $c\to+\infty$ in Cases A and B
The aim of this section is the study of the dynamics of (3.1) when $|c|\to+\infty$. Many results concerning the asymptotic behaviour of the solution of the FN model as $|c|\to+\infty$ are available in the literature – see [21, 22, 23]. Our concern is to match these results with the local bifurcation study performed in Section 3. We are particularly interested in Cases A and B whose results have been described in Theorems 3.4 and 3.9. By changing the time scale of (3.1), taking into account $\tau=t/c$ and $\varepsilon=1/c^{2},$ we get an equivalent form for (3.1) as follows (written with respect to the slow time $\tau$):
$\left\{\begin{array}[]{l}\varepsilon\dot{x}=\displaystyle\left[y-\left(\frac{x^{3}}{3}-x\right)\right]=:f(x,y,\varepsilon)\\ \\ \dot{y}=\displaystyle-(x+by-a)=:g(x,y,\varepsilon)\end{array}\right$. (4.1)
[…]
A fast-slow system consists of differential equations where the rates of change of some variables are much greater than those of others. This leads to a two time-scale system. The general approach to this type of system starts by grouping the variables in two disjoint sets: fast variable $x$ and slow variable $y$.
By decomposing the dynamics into slow and fast phases, each can be studied independently, allowing a clearer understanding of its underlying behaviour. The parameter $\varepsilon$ introduces this separation in system (4.1).
4.1. The critical manifold and bifurcations
We first analyse the singular case $\varepsilon=0.$ We follow the analysis performed in [10, 12, 13]. The critical manifold associated with (4.1) is the set
$\mathcal{C}_{0}=\{(x,y)\in\mathbb{R}^{2}:f(x,y,0)=0\}=\left\{(x,y)\in\mathbb{R}^{2}:y=\frac{x^{3}}{3}-x\right\}.$ (4.2)
Definition 4.1.
With respect to (4.1) with critical manifold $\mathcal{C}_{0}$:
  1. (1)
    The subset $S\subset\mathcal{C}_{0}$ is called normally hyperbolic provided that for all $e=(x_{0},y_{0})\in{S}$, we get:
    $\frac{\partial f}{\partial x}\left(e,0\right)\neq 0$
  2. (2)
    A normally hyperbolic set $S$ is called attracting (resp. repelling) when, for every $e\in{S}$, all eigenvalues of $\frac{\partial f}{\partial x}\left(e,0\right)$ possess negative (resp. positive) real part, respectively.
Definition 4.2.
With respect to (4.1) with critical manifold $\mathcal{C}_{0}$ (see (4.2)), the point $e\in\mathcal{C}_{0}$ is called a fold point if:
$\frac{\partial f}{\partial x}\left(e,0\right)=0\,,\quad\frac{\partial^{2}f}{\partial x^{2}}\left(e,0\right)\neq 0\quad\text{ and }\quad\frac{\partial f}{\partial y}\left(e,0\right)\neq 0.$
If $f\left(e,0\right)\neq 0$, the fold point is termed regular.
Coming back to system (4.1), we get $f\left(x,y,0\right)=x-\frac{x^{3}}{3}+y$ and, for $e=(x_{0},y_{0})\in\mathcal{C}_{0}$, one has:
$\frac{\partial f}{\partial x}\left(e,0\right)=1-x_{0}^{2}\neq 0\qquad\text{ if }\,\,x_{0}\neq\pm 1.$
Therefore, if $b=3/2$, the equilibria $E_{2}=\left(-1,2/3\right)$ and $E_{3}=(1,-2/3)$ are fold points (according to the Definition 4.2) since:
$\frac{\partial f}{\partial x}(E_{2,3},0)=0,\quad\frac{\partial^{2}f}{\partial^{2}x}(E_{2,3},0)\neq 0\quad\text{and}\quad\frac{\partial f}{\partial y}(E_{2,3},0)\neq 0$
and $S_{0}=\mathcal{C}_{0}\setminus\left\{E_{2},E_{3}\right\}$ is normally hyperbolic according to Definition 4.1. We may split $S_{0}$ into three disjoint open subsets:
$\mathcal{C}_{0\text{L}}$ $=\mathcal{C}_{0}\cap{\textstyle\left\{\left(x,y\right)\in\mathbb{R}^{2}:x<-1\right\}},$
$\mathcal{C}_{0\text{M}}$ $=\mathcal{C}_{0}\cap{\textstyle\left\{\left(x,y\right)\in\mathbb{R}^{2}:-1<x<1\right\}}\quad\text{and}$
$\mathcal{C}_{0\text{R}}$ $=\mathcal{C}_{0}\cap{\textstyle\left\{\left(x,y\right)\in\mathbb{R}^{2}:x>1\right\}}.$
Therefore, we get
$S_{0}=\mathcal{C}_{0\text{L}}\cup\mathcal{C}_{0\text{M}}\cup\mathcal{C}_{0\text{R}},$ (4.3)
where $\mathcal{C}_{0\text{L}}$ and $\mathcal{C}_{0\text{R}}$ are attracting and $\mathcal{C}_{0\text{M}}$ is repelling. As a result, trajectories of (4.1) in the fast flow move towards either to $\mathcal{C}_{0\text{L}}$ or $\mathcal{C}_{0\text{R}}$ and backwards to $\mathcal{C}_{0\text{M}}$ – see Figure 5.
Figure 5. (a) Illustration of the critical manifold associated to system (4.1). (b) Illustration of the dynamics of system (4.1) for the case $a=0,$ when there is a unique equilibrium.
Yet in the singular case ($\varepsilon=0$), depending on the values of $a,b\in\mathbb{R}$ in system (4.1), one, two, or three equilibria might arise as a consequence of the cubic curve’s intersection, $f(x,y,0)=0,$ and the line defined by xa+by=0x-a+by=0 ($\Leftrightarrow\dot{y}=0).$ If the system possesses three equilibria, only one of the following cases arise (cf. [10]):
Case 1:
each of the three equilibria lies in $\mathcal{C}_{0\text{M}}$ or else
Case 2:
each of the equilibria belongs to a different region $\mathcal{C}_{0\text{L}}$, $\mathcal{C}_{0\text{M}}$, and $\mathcal{C}_{0\text{R}}$.
When $\varepsilon=0,$ the independence of the time scales simplifies the system’s analysis. Fenichel’s Theorem [6] says that, near the normally hyperbolic part of $\mathcal{C}_{0}$, for $0<\varepsilon\ll 1,$ the perturbed system exhibits dynamics “analogous” to the singular system (in the sense of differentiable conjugacy), with a deviation of the order $\mathcal{O}\left(\varepsilon\right).$ Decreasing $\varepsilon$ makes the system trajectories increasingly similar to those of the singular system ($\varepsilon=0$). According to Fenichel’s Theorem, it is ensured that when $0<\varepsilon\ll 1,$ there is a set $S_{\varepsilon}$, $\mathcal{O}\left(\varepsilon\right)$–close to $S_{0}$ (in the $\mathbb{C}^{1}$–topology), which is locally hyperbolic and attracting (if $S_{0}$ is normally hyperbolic and attracting).
Definition 4.3.
The set $S_{\varepsilon}$ is known as the slow manifold for (4.1).
When the equilibrium of the system is a fold point, a subcritical (with respect to $b$) Hopf bifurcation arises, leading to the emergence of a periodic solution. In this section, the canard phenomenon described in [12, 13] links the local periodic solution generated by the Hopf bifurcation to the global limit cycle, like we proceed to explain (the canard point is correlated with the passage of a reduced problem equilibrium over the fold point under parameter variation – see Definition 4.6).
Definition 4.4.
A trajectory segment of a fast-slow system (4.1) is a canard if it stays within a $\mathcal{O}\left(\varepsilon\right)$–distance to a repelling branch of a slow manifold for a time that is $\mathcal{O}\left(1\right)$ on the slow time scale $\tau=t\varepsilon$.
Canards are related to equilibria of (4.1) that arise at fold points of the slow manifold. The next definitions deal with the general equation parametrised by $\lambda\in\mathbb{R}$:
$\left\{\begin{array}[]{l}\varepsilon\dot{x}=f(x,y,\lambda,\varepsilon)\\ \\ \,\dot{y}=g(x,y,\lambda,\varepsilon)\end{array}\right$. (4.4)
We now define a singular fold point, a different concept from that of a fold point; it is necessary to coincide with an equilibrium point of (4.1):
Definition 4.5.
With respect to (4.4), let $e=\left(x,y\right)$ denote a fold point. We refer to $e$ as a singular fold provided:
$f\left(e,\lambda,0\right)=0,$ $\frac{\partial f}{\partial x}\left(e,\lambda,0\right)=0,$ (4.5)
$\frac{\partial^{2}f}{\partial x^{2}}\left(e,\lambda,0\right)\neq 0,$ $\frac{\partial f}{\partial y}\left(e,\lambda,0\right)\neq 0\quad\text{and}\quad g\left(e,\lambda,0\right)=0.$
Definition 4.6.
A singular fold point $e$ (also known as a canard point) is called regular if:
$\frac{\partial g}{\partial x}\left(e,\lambda,0\right)\neq 0\quad\text{ and }\quad\frac{\partial g}{\partial\lambda}\left(e,\lambda,0\right)\neq 0.$ (4.6)
The dynamics near a regular singular fold point is explored in Theorem 4.7. The occurrence of canards is linked to a Hopf bifurcation, known in these systems as a singular Hopf bifurcation.
Theorem 4.7 ([12, 13], adapted).
Consider a fast-slow system of the form (4.4) where $\left(x,y\right)=\left(0,0\right)$ is canard point for $\lambda=0$, and
$\dot{x}=-y\,l_{1}\left(x,y,\lambda,\varepsilon\right)+x^{2}\,l_{2}\left(x,y,\lambda,\varepsilon\right)+\varepsilon\,l_{3}\left(x,y,\lambda,\varepsilon\right)$ (4.7)
$\dot{y}=\varepsilon\left(\pm x\,l_{4}\left(x,y,\lambda,\varepsilon\right)-\lambda\,l_{5}\left(x,y,\lambda,\varepsilon\right)+y\,l_{6}\left(x,y,\lambda,\varepsilon\right)\right)$
where
$l_{3}\left(x,y,\lambda,\varepsilon\right)=\mathcal{O}\left(x,y,\lambda,\varepsilon\right)\quad\text{and}\quad l_{j}\left(x,y,\lambda,\varepsilon\right)=1+\mathcal{O}\left(x,y,\lambda,\varepsilon\right),$
for $j\in\{1,2,4,5,6\}$.
Suppose that, when $\varepsilon=0$, a slow trajectory connects attracting and repelling regions of the critical manifold $\mathcal{C}_{0}$. Therefore, there is $\varepsilon_{0}>0$ and $\lambda_{0}>0$ so that for $0<\varepsilon<\varepsilon_{0}$ and $|\lambda|<\lambda_{0}$, the system possesses an equilibrium point $p\in\mathbb{R}^{2}$ around $(0,0)$ where $p$ approaches $(0,0)$ as $\left(\lambda,\varepsilon\right)$ tends to $(0,0)$. Moreover, there exists a smooth map $\lambda_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that assigns every $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a real value $\lambda$, leading to a family of canards, defined asymptotically as
$\lambda_{c}\left(\sqrt{\varepsilon}\right)=-\varepsilon\left(A+B\right)+\mathcal{O}\left(\varepsilon^{3/2}\right).$
and In addition, there is a continuous map $\lambda_{H}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each $\varepsilon\in\left[0,\varepsilon_{0}\right]$ to a value $\lambda$, associated to a Hopf bifurcations defined asymptotically as
$\lambda_{H}\left(\sqrt{\varepsilon}\right)=-\varepsilon B\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
where
$A=-\dfrac{\dfrac{\partial l_{1}}{\partial x}-3\dfrac{\partial l_{2}}{\partial x}+2\dfrac{\partial l_{4}}{\partial x}-2l_{6}}{8},\qquad B=\dfrac{\dfrac{\partial l_{3}}{\partial x}+l_{6}}{2},$
and the functions $l_{i},$ for $1\leq i\leq 6$, and their partial derivatives are calculated at $\left(x,y,\lambda,\varepsilon\right)=\left(0,0,0,0\right)$. The latter bifurcation is non-degenerate when $A\neq 0$, it occurs for positive values of $\lambda-\lambda_{H}$ if $A<0$, and for negative values of $\lambda-\lambda_{H}$ if $A>0$.
4.2. Case A $(a=0)$
[…]
$T_{F,-}^{2,3}(c)\leq T_{H}^{2,3}(c)\leq DH(c)\leq SNL(c).$ (4.8)
All the curves pass through $(b,c)=(1,1)$. Denote by $A_{1}\leq A_{2}<A_{3}<A_{4}$ the limits when $c\to+\infty$ of the previous curves (when it exists), respectively. More specifically, we have:
Figure 6. Partial scheme of the codimension 1 bifurcation curves emerging from the point $(b,c)=(1,1)$ associated with $f_{(0,b,c)}$ of (3.1), for $b,c>0$. The symbol $c^{\infty}$ represents “very large” values of $c$. The notation of the curves correspond to those of Theorem 3.4, for Case A. The line SNL is dashed because it does not make part of the DZ bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry. Near the curve $b=T_{H}^{2,3}$, the emerging periodic solutions give rise to canards as a consequence of Proposition 4.9.
Lemma 4.8.
As illustrated in Figure 6, the following equalities hold:
  1. (1)
    $A_{1}=\displaystyle\lim_{c\to+\infty}T_{F,-}^{2,3}(c)=3/2$
  2. (2)
    $A_{2}=\displaystyle\lim_{c\to+\infty}T_{H}^{2,3}(c)=3/2$
  3. (3)
    $A_{3}=\displaystyle\lim_{c\to+\infty}DH(c)={17/10}$
  4. (4)
    $A_{4}=\displaystyle\lim_{c\to+\infty}SNL(c)=+\infty$
In particular, we have $1<A_{1}\leq A_{2}<A_{3}<A_{4}.$
We omit the proof since it follows from straightforward computations using the expressions of Theorem 3.4 and Remark 3.6. A partial representation of these curves has been plotted in Figure 6.
Proposition 4.9.
With respect to (4.4), a smooth function $b_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ exists that assigns every $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a value $b$ leading to a family of canards in the system, asymptotically defined by:
$b_{c}\left(\sqrt{\varepsilon}\right)=\frac{3}{2}+\frac{5}{4}\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right).$
Proof.
The difference b3/2b-3/2 of system (4.4) is played by $\lambda$ of Theorem 4.7. The equilibrium $E_{3}$ is a singular fold point and a Hopf bifurcation occurs at $b=3/2$. Using Theorem 4.7, we perform a change of variables and parameter as
[…]
to obtain the equivalent system ($\tau=t/\varepsilon$)
$\left\{\begin{array}[]{l}\dot{\bar{x}}=\bar{y}-\bar{x}^{2}-\frac{1}{3}\bar{x}^{3}\\ \\ \dot{\bar{y}}=\varepsilon\left(-\bar{x}+\frac{2}{3}\lambda-\lambda\bar{y}-\frac{3}{2}\bar{y}\right).\end{array}\right$.
Using the notation introduced in Theorem 4.7, we obtain
$l_{1}=-1\qquad l_{2}=-1-\bar{x}/3\qquad l_{3}=0\qquad l_{4}=1\qquad l_{6}=-\frac{3}{2}$
hence A=1/2<0A=-1/2<0 and B=3/4<0B=-3/4<0 therefore the Hopf bifurcation occurs for values $\lambda>\lambda_{H}(\Leftrightarrow b>\frac{3}{2}+\lambda_{H})$ . Since for $b<T_{H}^{2,3}(c)$ the eigenvalues of $Df_{(0,b,c)},$ evaluated at $E_{i},$ $i=1,2,$ are negative, then the bifurcating periodic solution is unstable and canards occur when
$b_{c}(\sqrt{\varepsilon})=\frac{3}{2}+\frac{5}{4}\varepsilon+\mathcal{O}(\varepsilon^{3/2}).$
4.3. Digestive remark
We describe the dynamics of (3.1) in the case $a=0$, focusing on the first quadrant of the bifurcation diagram in the (b,c)(b,c)–plane. We suggest the reader follows the description by using Figure 6.
For $b\in(0,1)$ and $c>1$, there exists a unique equilibrium point $E_{1}$, which is unstable. This source is enclosed by a stable periodic solution denoted by $C^{s}$. At $b=1$, one of the eigenvalues of $Df_{(0,b,c)}$ at $E_{1}$ becomes zero, leading to the creation of two additional unstable nodes $E_{2}$ and $E_{3}$. In other words, $E_{1}$ undergoes a supercritical (with respect to $b$) Pitchfork bifurcation.
The saddles $E_{2}$ and $E_{3}$ subsequently become unstable foci and undergo a supercritical Hopf bifurcation, leading to two unstable periodic solutions (symmetric under $\kappa$). Around $b=DH(c)$, the invariant manifolds of $E_{1}$ evolve into a double homoclinic orbit, which is unstable, previously identified in [8].
As $b$ increases smoothly, the double homoclinic loop breaks apart, resulting in an unstable periodic orbit $C^{u}$, which surrounds both the stable and unstable manifolds of $E_{1}$. Near $b=SNL(c)$, the periodic orbits $C^{s}$ and $C^{u}$ collapse, and two stable equilibria remain, dominating the basin of attraction of any compact subset of $\mathbb{R}^{2}$. For $c\gg 1$, there is a small portion of the bifurcation parameter diagram $(b,c)$ of Figure 6, where the periodic orbits emerging from $T_{H}^{2,3}$ correspond to canards. They rapidly collapse into the double homoclinic cycle to $E_{1}$. Parameters $b_{c}$ that yield canards are only $\mathcal{O}\left(\varepsilon\right)$ away from the corresponding Hopf bifurcation values $b_{H}$ (see Figure 7):
$b_{c}-b_{H}=\frac{1}{2}\,\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right).$
Figure 7. Sketch of the singular bifurcations for $b\geq 1$ and $c<\infty$, in Case A. As $\varepsilon$ decreases, the interval $b_{c},b_{H}]$ becomes narrower. When $\varepsilon=0\,(\Leftrightarrow c=+\infty$), we have $b_{c}\equiv b_{H}$.
4.4. Case B $(b=0)$
Proposition 4.10.
With respect to (4.4), a continuous mapping $a_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ exists that assigns to each $\varepsilon\in\left[0,\varepsilon_{0}\right]$ a corresponding value $a$ giving rise to a canard, defined asymptotically by
$a_{c}\left(\sqrt{\varepsilon}\right)=-\frac{\varepsilon}{8}+\mathcal{O}\left(\varepsilon^{3/2}\right).$
Proof.
The difference a1a-1 of system (4.4) is played by $\lambda$ in Theorem 4.7. We are interested in only one of the two fold points of (3.1) located at (1,2/3)(1,-2/3). After shifting the fold point to the origin, reversing time $t\to-t$ and setting $\lambda=a-1$, we get
$\left\{\begin{array}[]{l}\dot{x}=-y+x^{2}(1+x/3),\\ \\ \dot{y}=\varepsilon(x-\lambda).\end{array}\right$.
In this standard form, the relevant parameters $\ell_{i}$ defined in Theorem 4.7 are
$\ell_{1}=-1,\quad\ell_{2}=1+x/3,\quad\ell_{3}=0,\quad\ell_{4}=1,\quad\ell_{5}=1,\quad\ell_{6}=0,\quad A=1.$
Hence, a maximal canard exists on a curve $\lambda_{c}$ in $(\varepsilon,\lambda)$–space defined by:
$a_{c}(\sqrt{\varepsilon})=-\frac{\varepsilon}{8}+\mathcal{O}(\varepsilon^{3/2}).$
4.5. Digestive remark
The dynamics of system (3.1) for $b=0$ is illustrated in Figures 4 and Table 2. Our analysis focuses on the first and fourth quadrants of the bifurcation diagram in the (a,c)(a,c)–plane and $c>0$. For a<1a<-1, the saddle $E_{1}$ becomes a stable node. Near a=1a=-1, a Belyakov transition occurs, followed by a supercritical Hopf bifurcation that gives rise to a stable periodic orbit that involves into a canard in an exponentially small portion of the bifurcation plane. This periodic solution disappears through another Hopf bifurcation at $a=1$.
5. The phase portraits “near infinity”
Studying the dynamics “near infinity” provides insight into the system’s asymptotic behaviour and allows the identification of trajectories that leave bounded sets. Using the techniques of Subsections 2.2 and 2.3, we compactify system (3.1) and describe the dynamics of (3.1) at infinity – this is what we call the dynamics “near infinity”. We distinguish between four cases
$b\geq 0,\,c>0,\qquad b\geq 0,\,c<0,\qquad b<0,\,c>0,\qquad b<0,\,c>0.$ (5.1)
The main theorem of this section is the following:
Theorem 5.1.
The dynamics of system (3.1) “near infinity” is topologically equivalent to one of the phase portraits depicted in Figure 8.
Insight into the reasoning
Before proceeding with the formal proof of the result, we give a geometrical interpretation of what is going to be done: we investigate the behaviour of trajectories of the planar vector field (3.1) as their position tends toward infinity. To do this, we apply the Poincaré compactification of Subsection 2.3, which transforms points at infinity into a bounded region represented by a disk containing a finite number of equilibria. We then study the qualitative dynamics of the “compactified” vector field paying particular attention to the equilibria located there, the directions along which trajectories diverge, and the conditions under which the flow at infinity becomes attracting or repelling. Some equilibria are singularities. In order to study their dynamics we need to desingularize them using the techniques of Subsection 2.2.
(a) $b\geq 0$ and $c>0$
(b) $b\geq 0$ and $c<0$
(c) $b<0$ and $c<0$
(d) $b<0$ and $c>0$
Figure 8. Dynamics near infinity on the Poincaré disc of system (3.1).
Proof.
Using (2.3) with $d=3$, system (3.1) in the local chart
$U_{1}=\{(y_{1},y_{2},y_{3})=(x_{1},x_{2},1)\in\mathbb{S}^{2}:y_{1}>0\}$
may be written as:
$\dot{u}=$ $-cu^{2}v^{2}-\frac{b}{c}uv^{2}-cuv^{2}+\frac{a}{c}v^{3}-\frac{1}{c}v^{2}+\frac{c}{3}u,$ (5.2)
$\dot{v}=$ $-cuv^{3}-cv^{3}+\frac{c}{3}v,$
where
$u=\frac{x_{2}}{x_{1}}\quad\text{and}\quad v=\frac{1}{x_{1}}.$
The system (5.2) has a unique equilibrium at the origin on $v=0$ (in any chart, points of $\mathbb{S}^{1}$ are characterized by $v=0$). The eigenvalues of the Jacobian matrix associated to (5.2) computed at the origin, are $\dfrac{c}{3},\dfrac{c}{3}.$ So, the origin is an unstable (resp. stable) node for $c>0$ (resp. $c<0$).
Analogously, using (2.4) and $d=3$, the system (3.1) in chart $U_{2}$ takes the form:
$\dot{u}=$ $\frac{1}{c}u^{2}v^{2}-\frac{a}{c}uv^{3}-\frac{c}{3}u^{3}+\frac{b}{c}uv^{2}+cuv^{2}+cv^{2},$ (5.3)
$\dot{v}=$ $\frac{1}{c}uv^{3}-\frac{a}{c}v^{4}+\frac{b}{c}v^{3}.$
The origin is an identically zero singular point for system (5.3). In order to desingularize this equilibrium, we are going to employ the quasi-homogeneous blow-up
$(u,v)=(\bar{u}\rho^{2},\bar{v}\rho^{3}),$ (5.4)
where $(\alpha,\beta)=(2,3)$ has been obtain by using the Newton diagram discussed in Subsection 2.2. The Newton polygon associated to (5.3) is illustrated in Figure 9.
Figure 9. The points/segments associated to (5.3) are $(2,2)$, $(1,3)$, $(3,0)$, $(1,2)$, $(0,2)$, $(0,4)$ and $(0,3)$. The Newton polygon is defined by the line $3y+2x\geq 6$. Therefore $(\alpha,\beta)=(2,3)$.
Applying the transformation (5.4) where $\bar{u}=1$ and $\bar{u}=-1,$ the $x$ directional blow-up, converts the system (5.3) into the vector fields defined by:
$X_{1}^{+}:\begin{cases}\frac{d{\rho}}{d\eta}=\frac{\rho\left(-3a\rho^{5}{\bar{v}}^{3}+3{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+3b{\rho}^{2}{\bar{v}}^{2}+3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{6c},\\ \\ \frac{d\bar{v}}{d\eta}=-\frac{\bar{v}\left(-a\rho^{5}{\bar{v}}^{3}+{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+b{\rho}^{2}{\bar{v}}^{2}+3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{2c},\end{cases}$
and
$X_{1}^{-}:\begin{cases}\frac{d{\rho}}{d\eta}=\frac{\rho\left(-3a\rho^{5}{\bar{v}}^{3}-3{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+3b{\rho}^{2}{\bar{v}}^{2}-3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{6c},\\ \\ \frac{d\bar{v}}{d\eta}=-\frac{\bar{v}\left(-a\rho^{5}{\bar{v}}^{3}+5{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+b{\rho}^{2}{\bar{v}}^{2}+3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{2c},\end{cases}$
respectively, where $d\eta=\rho^{4}d\tau$ (the old time variable is $\tau$ and the new one is $\eta$; when we calculate $X^{1}$, we can cancel $\rho^{4}$ by using this time rescaling).
We now need to analyze the equilibria at $\rho=0$ of $X_{1}^{\pm}.$ Regarding the vector field $X_{1}^{+},$ and $X_{1}^{-},$ one concludes that:
  1. Claim 1:
    The equilibria of system $X_{1}^{+}$ on $\rho=0$ are directly determined by the roots of
    $\bar{v}\left(3\bar{v}^{2}-1\right)=0.$
    There are three equilibria at $\rho=0$ denoted by
    $D_{1}:(0,0)\qquad\text{and}\qquad f^{\pm}:\left(0,\pm\dfrac{\sqrt{3}}{3}\right).$
    The following statements are valid:
    • 1.
      The equilibrium $D_{1}$ always is a saddle point for $X_{1}^{+}$.
    • 2.
      The points $f^{\pm}$ are semi-hyperbolic and the dynamics of $X_{1}^{\pm}$ restricted to the center manifold $\bar{v}=\mp\frac{\sqrt{3}(b+3c^{2})\rho^{2}}{18c^{2}}$ is given by $\dot{\rho}=\frac{b}{9c}{\rho^{3}}.$
      Proof.
      A center manifold in the form of $h(\rho):=\alpha_{1}\rho^{2}+h.o.t$ must satisfy the equation
      $\dot{\rho}\frac{\partial h}{\partial\rho}-\dot{\bar{v}}\big|_{(\rho,h(\rho))}=0.$
      After shifting $f^{\pm}$ to the origin, we obtain $\alpha_{1}=\mp\frac{\sqrt{3}(b+3c^{2})}{18c^{2}}.$ Then, the dynamics of $X_{1}^{\pm}$ restricted to the center manifold $\bar{v}=\mp\frac{\sqrt{3}(b+3c^{2})\rho^{2}}{18c^{2}}$ is given by $\dot{\rho}=\frac{b}{9c}{\rho^{3}}$ (more details in [5, Theorem 2.19]). ∎
    • 3.
      The equilibria $f^{\pm}$ are:
      • 1a.
[…]
      • 2a.
        an unstable node for $\{(a,b,c)\big|b<0,\,c<0\}$ – see Figure 10(c).
      • 3a.
        a stable node for $\{(a,b,c)\big|b<0,\,c>0\}.$ – see Figure 10(d).
  2. Claim 2:
    The equilibria of system $X_{1}^{-}$ on $\rho=0$ are determined by the roots of
    $\bar{v}\left(3\bar{v}^{2}-1\right)=0.$
    There are three equilibria at $\rho=0$ denoted by $D_{2}:(0,0)$ and $g^{\pm}:\left(0,\pm\frac{\sqrt{3}}{3}\right),$ and the results of Claim 1 are valid in this case (changing $D_{1}$ and $f^{\pm}$ by $D_{2}$ and $g_{\pm}$, respectively).
    Using the transformation (5.4) with $\bar{v}=1$ and $\bar{v}=-1,$ the $y$ directional blow-up transforms system (5.3) into
    $Y_{1}^{+}:\begin{cases}\frac{d{\rho}}{d\eta}=\mathcal{O}(|\rho,\bar{u}|),\\ \\ \frac{d\bar{u}}{d\eta}=c,\end{cases}\quad\quad Y_{1}^{-}:\begin{cases}\frac{d{\rho}}{d\eta}=\mathcal{O}(|\rho,\bar{u}|),\\ \\ \frac{d\bar{u}}{d\eta}=c,\end{cases}$
    respectively, where $d\eta=\rho^{4}d\tau.$ We now need to analyze the point (0, 0) of $Y_{1}^{\pm}.$ Indeed,
  3. Claim 3:
    For the vector fields $Y_{1}^{\pm},$ point (0, 0) is a regular point (not an equilibrium).
From the Claims 1–3, we know the dynamics of the vector fields $X_{1}^{\pm}$ and $Y_{1}^{\pm}$. For the vector fields $Y_{1}^{\pm},$ $(0,0)$ is a regular point. Thus, Figure 10 illustrates the qualitative properties of the vector fields $X_{1}^{\pm}$ and $Y_{1}^{\pm}.$ By shrinking the circles in these figures to the origin (as the radii of the circles decrease to zero and the circles collapse at $(0,0)$ ), we achieve to the phase portraits shown in Figure 11. So, as shown in Figure 8, the equilibrium of chart $U_{2}$ is a saddle with four hyperbolic sectors for $b\geq 0.$ Moreover, it becomes an unstable (resp. stable) node when $b,c<0$ and $b<0,c>0,$ respectively.
(a) $b\geq 0$ and $c>0$
(b) $b\geq 0$ and $c<0$
(c) $b<0$ and $c<0$
(d) $b<0$ and $c>0$
Figure 10. Qualitative properties of vector fields $X_{1}^{\pm}$ and $Y_{1}^{\pm}.$ The axes corresponding to systems $X_{1}^{\pm}$ and $Y_{1}^{\pm}$ are $(\rho,\bar{v})$ and $(\rho,\bar{u}),$ respectively.
(a) $b\geq 0$ and $c>0$
(b) $b\geq 0$ and $c<0$
(c) $b<0$ and $c<0$
(d) $b<0$ and $c>0$
Figure 11. Local phase portraits of system (5.3) (the representation of system (3.1) in the chart $U_{2}$) at $(0,0)$.
6. Discussion and concluding remark
Although the FHN equations were originally developed as a simplified representation of nerve impulses, they have also been extensively investigated in purely mathematical purposes [12, 13, 10, 23] – they consist of a very simple model of equations with rich dynamics.
Results obtained on [21, pp. 180] synthesised the global bifurcation diagram for the FHN system (3.1). It has been obtained by putting together, as in a “huge puzzle”, all local bifurcation diagrams obtained. The authors have concentrated their attention to a particular parameter region relevant to physiology ($|c|>1+\sqrt{3}$).
Trying to complete and understand the bifurcation diagram of (3.1), in this paper we have discussed the finite equilibria of (3.1), as well as their bifurcations, for the three scenarios identified in Subsection 1.4:
Case A:
$a=0,$
Case B:
$b=0,$
Case C:
$a\neq 0$ and $0<b<1$.
The richest scenario is Case A. In Theorem 3.4 we have found a Double-zero Bifurcation with a $\mathbb{Z}_{2}(\kappa)$–symmetry. We have obtained precise expressions of the bifurcation curves passing through the bifurcation points $(b,c)=(1,\pm 1)$, complementing the work started in [8]. We also give an analytical proof of the results stated in Section 2.4 of [20].
We have been able to explain rigorously the dynamics of Regions 8, 11, 17 and 18 of [21] on the line defined by $a=0$ and $b>0$, as we proceed to explain:
  • Point Q: Pitchfork bifurcation of $E_{1}$ (Theorem 3.4, curve $b=T_{P}$);
  • Region 8: existence of three equilibria, one saddle $E_{1}$ and two sources $E_{2},E_{3}$;
  • Point $\textbf{Q}_{0}$: Hopf bifurcations of $E_{2},E_{3}$ (Theorem 3.4, curve $b=T_{H}^{2,3})$;
  • Region 11: three periodic solutions (one stable, two unstable);
  • Point $\textbf{Q}_{6}$: Double Homoclinic (Theorem 3.4, curve b=DHb={DH});
  • Region 17: two periodic solutions of different stabilities;
  • Point T: Saddle-node bifurcation of non-hyperbolic solutions, making part of the Bautin bifurcation of $\textbf{Q}_{17}$ and $\textbf{Q}_{18}$ – see Remark 3.6;
  • Region 18: three equilibria $E_{1},E_{2}$ and $E_{3}$.
The precise location the above points/regions depend on $c$ but their relative position does not.
The non-hyperbolic equilibria associated with parameter values located at Q (see [21]) are degenerated saddle-node (cusp) equilibria, which are attracting for $c<1$ and repelling for $c>1$. For $c=1$, following the Remark 3.6, this point might correspond to a degenerated Bogdanov-Takens of order two (with symmetry) – a kind of codimension-three bifurcation. As pointed out in [23], this would be the unique generic codimension-three local bifurcation exhibited by the FHN model (3.1) and its complete understanding is an open problem. Using the same line of argument, in Cases B and C, we have analysed the dynamics of (3.1).
In Section 4, we have studied the asymptotic case $c\to\pm\infty$, where canards are detected. A key point in this research is the empirical difficulty to detect canards since the smaller value of $\varepsilon$ is, the narrower the interval of $b$ (Case A) or $a$ (Case B) values for which canards appear – see Figure 7. The dynamics that emerges in the system within these narrow parameter ranges is known as the canard explosion and has been studied in [10, 14, 12].
In Section 5, we have continued the analysis with the compactification of the phase portraits of (3.1) on the Poincaré disc. This brings additional information of the trajectories which come from or tend to infinity. We have provided phase and bifurcation diagrams in all the three cases.
Throughout this article, we have connected the local bifurcation theory (Theorems 3.4, 3.9 and Proposition 3.10) with dynamics “near infinity” (Theorem 5.1). Our contributions do not finish the whole discussion of the bifurcation analysis of (3.1); bifurcations might make part of high codimension phenomena.
Theoretically, the maximum number of non-trivial periodic solutions of (3.1) is three [21, pp. 218]. There are regions where the dynamics is completely determined and the phase portrait is complete and others in which the complete analysis of the bifurcation diagram of [21, pp. 180] is still ongoing.
7. Appendix
All phase portrait are illustrated with numerical simulations in Tables 1 and 2 using the software pplane9. Initial conditions were rigorously chosen to illustrate the characteristic behaviour in each region. Furthermore, limit cycles are indicated by the red trajectories in the phase plot. The phase portraits of Table 1 and 2 have been plotted for the parameter values,
$(c,b)=$ $(2.5,1.1),\;(2,1.2),\;(2,1.3),\;(2,1.34),\;(1.3,1.3),\;(0.5,1.1),\;(-0.5,1.1),\;(-2,1.5),$
$(-2,1.34),\;(-2,1.33),\;(-2,1.2),\;(-2.5,1.1),\;(-3,0.5),\;(-1,0.5),\;(-0.5,0.4)\;$
$(-0.1,0.9),\;(0.1,0.9),\;(0.4,0.4),\;(1.5,0.5),\;(3,0.5),\;(2.5,-0.5),\;(2,-0.5),\;$
$(1,-1),\;(0.2,-1),\;(-0.2,-1),\;(-1,-1),\;(-1.5,-0.5),\;(-2.5,-0.5),$
and
$(c,a)=$ $(1,0.2),\;(-1,0.2),\;(5,2),\;(-5,2),\;(-1,0.5)\;(1,0.5)\;(-5,0.5)\;(5,0.5),$
respectively.
[Uncaptioned image] Region 1 [Uncaptioned image] Region 2 [Uncaptioned image] Region 3 [Uncaptioned image] Region 4 [Uncaptioned image] Region 5 [Uncaptioned image] Region 6 [Uncaptioned image] Region 7
[Uncaptioned image] Region 8 [Uncaptioned image] Region 9 [Uncaptioned image] Region 10 [Uncaptioned image] Region 11 [Uncaptioned image] Region 12 [Uncaptioned image] Region 13 [Uncaptioned image] Region 14
[Uncaptioned image] Region 15 [Uncaptioned image] Region 16 [Uncaptioned image] Region 17 [Uncaptioned image] Region 18 [Uncaptioned image] Region 19 [Uncaptioned image] Region 20     [Uncaptioned image] Region 21
[Uncaptioned image] Region 22 [Uncaptioned image] Region 23 [Uncaptioned image] Region 24 [Uncaptioned image] Region 25 [Uncaptioned image] Region 26 [Uncaptioned image] Region 27 [Uncaptioned image] Region 28
Table 1. The local phase portraits corresponding to Case A, which is associated with system (3.1) when $a=0,$ for Regions 1–28 of Figure 3.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Region 1 Region 2 Region 3 Region 4
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Region 5 Region 6 Region 7 Region 8
Table 2. The local phase portraits corresponding to Case B, which is associated with system (3.1) when $b=0,$ for Regions 1–8 of Figure 4.